An analyst has compiled the following information regarding Rubsam, Inc., a calendar year fiscal year company:

Liabilities 12/31                      $1,000

Paid in Capital 12/31                 $500

Retained Earnings 1/1               $600

Revenue 1/1-12/31                $5,000

Expenses 1/1-12/31                $4,300

There have been no distributions to stockholders. The most likely estimate of Rubsam’s total assets at 12/31 should be closest to:

  1. $2,100
  2. $2,300
  3. $2,800
Answer #1

As we know

Total assets = Total Liabilities + Stockholder's Equity

Ending retained earnings = Beginning retained earnings + Net income

= 600 + (5000-4300) = 600 + 700 = 1300

Total assets = 1000 + 500 + 1300

= 2800

Option C